OpenStudy (anonymous):

How would I solve this: lim(x->0+)(sin((pi/2)e^sqrt(x)))

OpenStudy (anonymous):

Never mind. I figured it out. I feel kind of silly now.

OpenStudy (anonymous):

$$\lim_{x\to0^+}\sin\left(\frac\pi2e^\sqrt{x}\right)=\sin\left(\frac\pi2e^0\right)=1$$

OpenStudy (anonymous):

Yep. That's what I got. I was trying to overcomplicate it too much.

OpenStudy (anonymous):

It's a right-hand limit because the function is only defined on \([0,\infty)\); if you recall the definition of a limit, we require our point to be in the interior of its domain (not the boundary) so that we can 'approach' from both sides.